Circular No. 8239

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 W1 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below) was posted on the NEO Confirmation Page and has been found to show cometary appearance by several CCD observers, including J. Young (Table Mountain, CA, 0.6-m reflector; Nov. 17.1 UT, very round coma of diameter 8′′ and mag 17.5 with a faint, featureless tail 50′′ long in p.a. 42^{o}), R. Fredrick and R. Trentman (Louisburg, KS, 0.75-m reflector; Nov. 17.1, very diffuse tail ≈ 4′′ long in p.a. 20^{o}), J. E. Rogers (Camarillo, CA, 0.30-m reflector; Nov. 17.1, diffuse), and J. Lacruz (Madrid, Spain, 0.30-m reflector; Nov. 17.8, diffuse coma extending some 50′′ to the north).

2003 UT  α_{2000}  δ_{2000}  Mag.
Nov. 16.08352  17^{h}19^{m}24.^{s}46  +61^{o}31′38′′.2  18.5

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-W04.

T = 2003 Nov. 14.185 TT  ω = 116.^{o}350
Ω = 256.288 }2000.0
q = 1.65707 AU  i = 78.342

COMET 2P/ENCKE L. M. Woodney, University of Central Florida; and T. C. Owen and

Y. R. Fernandez, University of Hawaii, report the detection of HCN from comet 2P/Encke. The HCN J(4–3) transition was observed during Nov 9– 11 UT at the James Clerk Maxwell Telescope. The line had a FWHM of 1.4 km/s and an integrated line strength of 0.057 ± 0.011 K km/s. Assuming a rotational temperature of 43 K, and using a Haser model, a production rate of Q(HCN) = 9.8 × 10^{23} molecules/s was derived.

Visual total-magnitude and coma-diameter estimates: Oct. 24.85 UT, 9.8, 8′ (E. P. Bus, Groningen, The Netherlands, 0.08-m refractor); Nov. 9.05, 8.5, 14′ (B. H. Granslo, Lygna, Norway, 0.15-m reflector; during total lunar eclipse); 11.68, 8.4, 12′ (P. Guzik, Krosno, Poland, 10×50 binoculars); 12.71, 8.2, 15′ (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 14.73, 7.8, 15′ (T. Scarmato, Calabria, Italy, 7×50 binoculars); 16.06, 7.7, ∼ 9′ (D. W. E. Green, Hubbardston, MA, 12×50 binoculars).

2003 November 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green